Perform the indicated operation:
a) ((x^2+x-2)/x )/ (x+2)/x^2
b) (1/x-3)+((x-1)/(x^2+3x+9))+(x^2+x-3)/(x^3-27)

Answer by maali40(13)

((x^2+x-2)/x )/ (x+2)/x^2 is(x-1)*(x+2)/x/(x+2)/x^2=(x-1)/x^3 unless x=-2 then it's 1/x^2
(1/x-3)+((x-1)/(x^2+3x+9))+(x^2+x-3)/(x^3-27)
is ((x-1)*(x-3)+x^2+x-3)/(x^3-27)+(1-3x)/x unless x=3 then it's (0-8)/3+(2+0)/27+(9+3*0)/27/0 where 0 -+1+-1 " -1+1 or 1-1"